OpenStudy (anonymous):

Definite integral gives you the area, and area can't be negative. But while doing exercises, i cam across a lot of definite integrals, which after integrating, were giving negative answers. Why is it so?

OpenStudy (amistre64):

Area in a direction can be negative; maybe

OpenStudy (amistre64):

its either that or youve got your numbers backwards and are subtracting the smaller from the larger

OpenStudy (amistre64):

15 - 3 = 12 3 - 15 = -12

OpenStudy (angela210793):

well sometimes the area equals 0...so...that may b true...

OpenStudy (anonymous):

For example, when you integrate this definite integral, \[\int\limits_{-1}^{-3} (x-1)^{-3} dx\] the answer is -3/32 is it so because of the negative limits?

OpenStudy (anonymous):

hullo?

OpenStudy (amistre64):

You are simply on the "other side" of normal

OpenStudy (anonymous):

Ummm?

OpenStudy (amistre64):

you can flip your bounds and it makes a (-) in front

OpenStudy (amistre64):

it simply means that the area that you want is to the left of x=0

OpenStudy (amistre64):

since you are working in a region that is "negative" to begin with, it gonne be a "negative" area

OpenStudy (anonymous):

Alright, thank you...

OpenStudy (anonymous):

Are you a teacher in some college?

OpenStudy (amistre64):

nope...

OpenStudy (anonymous):

You're a student? Whoa, man than your genius.

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

then **

OpenStudy (anonymous):

No, really, are you a student?

OpenStudy (amistre64):

I am a college student yes; 35 years old tho so that might account for some wits lol

OpenStudy (anonymous):

of which year then?

OpenStudy (anonymous):

then you must be a Master's student, Well anyways thank you. Take care.

OpenStudy (anonymous):

Well, I'm a 12th year student, just to tell you, by the way.

OpenStudy (amistre64):

:) youll grow into it